### v2

# A Multi-Agent Approach to Fault Localization via Graph-Based Retrieval and Reflexion

Abstract.
Identifying and resolving software faults remains a challenging and resource-intensive process. Traditional fault localization techniques, such as Spectrum-Based Fault Localization (SBFL), leverage statistical analysis of test coverage but often suffer from limited accuracy. While learning-based approaches improve fault localization, they demand extensive training datasets and high computational resources. Recent advances in Large Language Models (LLMs) offer new opportunities by enhancing code understanding and reasoning. However, existing LLM-based fault localization techniques face significant challenges, including token limitations, performance degradation with long inputs, and scalability issues in complex software systems. To overcome these obstacles, we propose LLM4FL, a multi-agent fault localization framework that utilizes three specialized LLM agents. First, the Context Extraction Agent applies an order-sensitive segmentation strategy to partition large coverage data within the LLM’s token limit, analyze failure context, and prioritize failure-related methods. The extracted data is then processed by the Debugger Agent, which employs graph-based retrieval-augmented code navigation to reason about failure causes and rank suspicious methods. Finally, the Reviewer Agent re-evaluates the identified faulty methods using verbal reinforcement learning, engaging in self-criticism and iterative refinement. Evaluated on the Defects4J (V2.0.0) benchmark, which includes 675 faults from 14 Java projects, LLM4FL achieves an 18.55% improvement in Top-1 accuracy over AutoFL and 4.82% over SoapFL. It outperforms supervised techniques such as DeepFL and Grace, all without requiring task-specific training. Furthermore, its coverage segmentation and prompt chaining strategies enhance performance, increasing Top-1 accuracy by up to 22%.
Fault localization, large language model, agent
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
The process of locating and fixing software faults requires significant time and effort. Research shows that software development teams allocate more than half of their budgets to testing and debugging activities (Hait and Tassey, 2002; Alaboudi and LaToza, 2021). As software systems become increasingly complex, the demand for more accurate fault localization techniques grows. Hence, to assist developers and reduce debugging costs, researchers have developed various fault localization techniques (Li et al., 2019; Lou et al., 2021; Qian et al., 2023b; Li et al., 2021; Abreu et al., 2009; Sohn and Yoo, 2017). These techniques analyze code coverage and program execution to identify the most likely faulty code, assisting developers in finding the fault.
Despite the advances in fault localization, many existing techniques still struggle with scalability and precision. Traditional methods, such as Spectrum-Based Fault Localization (SBFL), use statistical analysis to analyze coverage data from passing and failing test cases to rank suspicious code elements (Abreu et al., 2006). While these techniques provide valuable insights, their accuracy is lower. Their reliance on statistical correlations between test failures and code coverage does not always capture the deeper semantic relationships needed for more accurate fault localization (Wong et al., 2016; Xie et al., 2013; Le et al., 2013). Recent techniques applied machine learning (Sohn and Yoo, 2017; Zhang et al., 2019b; Li et al., 2021, 2019) and deep learning models (Lou et al., 2021; Qian et al., 2023b; Rafi et al., 2024; Lou et al., 2021) to address these issues to improve fault localization. These methods enhance the ranking of suspicious code elements by incorporating additional information like code complexity, text similarity, and historical fault data. However, these techniques often require extensive training data that may not be available.
Recent advances in Large Language Models (LLMs) have shown great potential for software fault localization due to their strong language comprehension and generation capabilities (Abedu et al., 2024; Lin et al., 2024b). LLMs trained on extensive programming datasets can understand code structure, interpret error messages, and even suggest bug fixes (Wu et al., 2023; Pu et al., 2023; Li et al., 2023). These models, with their ability to analyze and process both natural language and code, present an opportunity to significantly improve traditional fault localization methods by incorporating deeper semantic analysis and context-aware reasoning.
However, there are several challenges in using LLMs for fault localization. Firstly, LLMs face significant difficulties when analyzing large-scale repository-level data, primarily due to inherent token limitations, as highlighted in recent studies (Hadi et al., 2023; Hou et al., 2023; Wu et al., 2023). Secondly, the performance of LLMs tends to decrease when they are applied to complex systems that require reasoning about multiple interacting components. This complexity can negatively impact the performance of LLMs (Liu et al., 2024; Wu et al., 2023). LLMs may struggle to understand the inter-procedural method call relationships, an essential aspect for understanding how faults propagate across the entire software repository (Chatterjee et al., 2015; Levy et al., 2024; Liu et al., 2024).
In this paper, we propose LLM4FL, a multi-agent LLM-based fault localization technique designed to analyze large-scale software projects at the repository level. LLM4FL leverages a multi-agent framework that emulates developers’ fault localization process. LLM4FL employs three LLM agents, each with a specialized role to enhance fault localization effectiveness. The Context Extraction Agent divides the methods covered by the failing test cases into smaller groups, each within LLM’s token limitation. Then, the agent analyzes the test code and test failure messages to reason about possible failure causes. This helps prioritize the most suspicious covered methods from each group before passing them on to the next agent. Next, the Debugger Agent analyzes each covered method and performs graph-based, retrieval-augmented (Graph-RAG) code navigation. It traverses the prioritized methods’ code iteratively along call dependencies to analyze and rank suspicious methods. Finally, the Reviewer Agent re-ranks the identified fault methods collected from all groups using verbal reinforcement learning (Shinn et al., 2024), criticizing and refining each response to ensure more precise fault localization.
We evaluated LLM4FL using the Defects4J (V2.0.0) benchmark (Just et al., 2014), which contains 675 real-world faults from 14 open-source Java projects. Our results demonstrate that LLM4FL surpasses LLM-based technique AutoFL (Kang et al., 2024) and AgentFL/SoapFL (Qin et al., 2024) by achieving 18.55% and 4.82% higher Top-1 accuracy, respectively. LLM4FL is also the cheapest among the three, costing only \$0.05 per fault. Additionally, LLM4FL outperforms supervised techniques such as DeepFL (Li et al., 2019) and Grace (Lou et al., 2021), even without task-specific training. We also analyzed the impact of individual components within LLM4FL on fault localization accuracy. Our findings indicate that each component plays a significant role in performance. Among these, dividing the covered methods into smaller groups, and Graph-RAG-based code navigation contributes the most. Moreover, we examined whether the initial ordering of methods provided to the LLM influences performance. The results reveal that method ordering is important. Even though LLM eventually visits all methods, the Top-1 accuracy still has a difference of up to 22% when comparing an execution-based ordering and the order provided by DepGraph (Rafi et al., 2024).
The paper makes the following contributions:
  • We introduce LLM4FL, a novel LLM-based fault localization technique that employs a divide-and-conquer strategy. This technique groups large coverage data and ranks the covered methods using an SBFL formula. Using multiple agents and code navigation based on Graph-RAG, LLM4FL analyzes the repository iteratively to identify and localize faults.
  • We conducted an extensive evaluation and LLM4FL demonstrates superior performance, surpassing AutoFL (Kang et al., 2024) by 18.55% and AgentFL/SoapFL (Qin et al., 2024) by 4.82% in Top-1 accuracy. It also outperforms supervised techniques like DeepFL and Grace, achieving these results without requiring task-specific training.
  • Our analysis of LLM4FL’s components shows that key components like dividing methods into smaller groups and code navigation are essential to its fault localization accuracy. Removing these features leads to performance declines, emphasizing their importance in handling token limitations and effective fault analysis.
  • We find that the different ordering on the initial method list passed to LLM can affect fault localization accuracy by up to 22% in Top-1 scores, even though LLM eventually visits all the methods.
In short, we provide a strategy to mitigate the token limitation issues and analyze repository-level data in LLM-based fault localization. We also highlight the impact of initial method ordering for LLM’s input. The findings may help inspire future research on LLM-based fault localization for large-scale software projects.
Paper Organization. Section 2 discusses background and related work. Section 3 describes our technique, LLM4FL. Section 4 presents the experimental results. Section 5 discusses the threats to validity. Section 6 concludes the paper.
Table 1. Comparison of LLM-based fault localization techniques.
Aspect LLM4FL (this work) AutoFL (Kang et al., 2024) SoapFL/AgentFL (Qin et al., 2024) AutoCodeRover (Zhang et al., 2024)
Navigating large code repository The agent autonomously navigates the repository, iteratively following an inter-procedural call graph to retrieve and explore connected methods. Retrieve class and method signatures related to the failure, then let the LLM decide which methods need deeper analysis. Analyze the test failures to identify the most suspicious classes, then iteratively analyze related methods. Start with the file or method names extracted from issue reports, then iteratively search the repository based on called methods in the retrieved source code.
Handling token limitation Divides methods into groups and analyzes each group separately. Provide only class and method signatures to the LLM, then let it decide which methods need deeper analysis. Use LLMs to generate description for classes and methods using their signatures. Then, the LLM decides which classes/methods are relevant to the test failures for further analysis. Iteratively search for relevant classes and methods based on the issue description and prioritize methods ranked by SBFL techniques.
Ranking of faulty methods Hierarchical ranking. First, it finds the suspicious methods and ranks them in each group. Then, it employs verbal reinforcement learning to combine and re-rank the suspicious methods from all groups. Re-run the fault localization process multiple times and do a majority vote on the result. Utilize a multi-round dialogue strategy that allows LLM agents to discuss the code to score and rank faulty methods. Does not rank methods; it selects a method as faulty from the retrieved context and generates a patch.
2. Background and Related Work
2.1. Background
Large Language Models. Large Language Models (LLMs), primarily built on the transformer architecture (Meta AI, 2024; Brown, 2020; Roziere et al., 2023), have significantly advanced the field of natural language processing (NLP). These LLMs, such as the widely recognized GPT3 model with its 175 billion parameters (Brown, 2020), are trained on diverse text data from various sources, including source code. The training involves self-supervised learning objectives that enable these models to develop a deep understanding of language and generate contextually relevant and semantically coherent text. LLMs have shown substantial capability in tasks that involve complex language comprehension and generation (Abedu et al., 2024; Lin et al., 2024b), such as code recognition and generation. Recent research has leveraged LLMs in software engineering tasks, particularly in fault localization (Kang et al., 2024; Qin et al., 2024; Yang et al., 2024), where they assist in identifying the faulty code groups responsible for software errors. One of the key advantages of using LLMs in fault localization is their ability to process both natural language and code without re-training, allowing them to analyze error messages, stack traces, and test case information to infer suspicious methods or code sections in an unsupervised zero-shot setting.
LLM Agents. LLM agents leverage LLMs to autonomously execute tasks described in natural language, making them versatile tools across various domains. LLM agents are artificial intelligence systems that utilize LLMs as their core computational engines to understand questions and generate human-like responses. They leverage functionalities like memory management (Zhou et al., 2023) and tool integration (Xia et al., 2024; Roy et al., 2024) to handle multi-step and complex operations seamlessly. The agents can refine their responses based on feedback, learn from new information, and even interact with other AI agents to collaboratively solve complex tasks (Hong et al., 2024; Qian et al., 2023a; Xu et al., 2023; Lin et al., 2024b). Through prompting, agents can be assigned different roles (e.g., a developer or a tester), providing more domain-specific responses that help improve the answer (Hong et al., 2024; White et al., 2024; Shao et al., 2023).
Recent studies (Shinn et al., 2024; Renze and Guven, 2024; Pan et al., 2025) explore using the agent’s verbal reasoning result to guide iterative improvement (i.e., verbal reinforcement learning), which has shown promising improvement in downstream tasks. In verbal reinforcement learning, LLM agents receive natural language feedback, such as reasoning results or instructions, from other agents as a reward signal. This allows agents to learn and adapt their behavior based on human-like guidance, improving their learning process to solve complex tasks. As the capabilities of large language model (LLM) agents grow, they play an essential role in enhancing automation and increasing productivity in software development. They can assist in code generation (Nijkamp et al., 2022; Lin et al., 2024a; Gu, 2023), debugging (Lee et al., 2024; Kang et al., 2024), test case creation (Huang et al., 2024; Chen et al., 2024b), and automated refactoring (Pomian et al., 2024; Liu et al., 2025), enabling developers to streamline repetitive tasks and focus on higher-level design and problem-solving. Additionally, LLM agents can facilitate collaborative software engineering by acting as intelligent assistants in code reviews, documentation generation, and issue resolution, improving overall development efficiency (Lin et al., 2024a; He et al., 2024). In this paper, we explore using LLM agents to improve fault localization by emulating developers’ debugging process using verbal reinforcement learning.
2.2. Related Work
Spectrum-based Fault Localization. Spectrum-based fault localization (SBFL) (Abreu et al., 2006; Jones et al., 2002; Wong et al., 2013; Abreu et al., 2009) employs statistical techniques to evaluate the suspiciousness of individual code elements, such as methods, by analyzing test outcomes and execution traces. The core idea of SBFL is that code components that are executed more frequently in failing tests and less frequently in passing tests are more likely to contain faults. Despite its widespread study, SBFL’s practical effectiveness remains limited (Kochhar et al., 2016; Xie et al., 2016). To enhance SBFL’s accuracy, recent research (Cui et al., 2020; Wen et al., 2019; Chen et al., 2022; Xu et al., 2020b) has suggested incorporating additional data, such as code changes (Wen et al., 2019; Chen et al., 2022) or mutation analysis (Cui et al., 2020; Xu et al., 2020b). However, SBFL’s reliance on code coverage metrics still poses challenges, as its suspiciousness scores may not generalize effectively to different faults or systems.
Learning-based fault localization. Recent efforts have focused on improving SBFL with learning-based methods (Sohn and Yoo, 2017; Zhang et al., 2019b; Li et al., 2021; Li and Zhang, 2017; Li et al., 2019; Zhang et al., 2019a). These approaches use machine learning models like radial basis function networks (Wong et al., 2011), back-propagation networks (Wong and Qi, 2009), and convolutional neural networks (Zhang et al., 2019a; Li et al., 2021; Albawi et al., 2017) to estimate suspiciousness scores based on historical faults. Some techniques, such as FLUCCS (Sohn and Yoo, 2017), combine SBFL scores with metrics like code complexity, while others, like DeepFL (Li et al., 2019) and CombineFL (Zou et al., 2019), merge multiple sources such as spectrum-based and mutation-based data (Moon et al., 2014; Papadakis and Le Traon, 2015; Dutta and Godboley, 2021). Graph neural networks (GNNs) have also been applied to fault localization (Qian et al., 2023b; Lou et al., 2021; Qian et al., 2021; Xu et al., 2020a). Techniques like Grace (Lou et al., 2021) and GNET4FL (Qian et al., 2023b) utilize test coverage and source code structure for improved accuracy, while DepGraph (Rafi et al., 2024) refines these approaches by graph pruning and incorporating code change information, resulting in higher performance with reduced computational demands. Although these learning-based techniques show improved results, they require training data that may not be available to every project.
LLM-Based Fault Localization. Large Language Models (LLMs), such as GPT-4o (OpenAI, 2024), LLaMA (Meta AI, 2024), and ChatGPT (Achiam et al., 2023), demonstrated remarkable abilities in processing both natural and programming languages. LLMs have shown potential in identifying and fixing errors using program code and error logs (Achiam et al., 2023). Some LLM-based fault localization techniques operate on small code snippets due to LLM’s token limitations. LLMAO (Yang et al., 2024) uses bidirectional adapters to score suspicious lines within a 128-line context. In contrast, Wu et al. (Wu et al., 2023) prompt ChatGPT with code and error logs but struggle to scale to large projects (Liu et al., 2024).
Recent fault localization techniques employ different strategies to navigate large codebases and refine fault localization results (Kang et al., 2024; Zhang et al., 2024; Qin et al., 2024). Each technique consists of key components for retrieving failure-related classes or methods, navigating the code repository, and ranking suspicious methods. However, there are differences in how these techniques design the components. Table 1 provides a summary of the key differences between prior techniques and ours. AutoFL (Kang et al., 2024) retrieves class and method signatures related to the test failure, and prompts the LLM to decide which methods need future analysis (i.e., to examine the code). While this approach ensures the input size is less than the LLM’s token limitation, it does not consider method dependence in the analysis. Instead, it refines rankings by repeatedly running the fault localization process and using a majority voting mechanism to determine the most likely faulty methods. AgentFL/SoapFL (Qin et al., 2024) models fault localization as a structured operating procedure with controlled phases. It uses a document-guided approach, where the system creates (or refines) LLM-generated or developer-provided documentation for classes and methods, then navigates the codebase based on these documents. By iteratively scoring and ranking suspicious methods, it continuously narrows down the fault candidates. It integrates document-guided search and enhanced method documentation to navigate the codebase while using iterative LLM-based scoring and ranking to refine the set of suspicious methods. AutoCodeRover (Zhang et al., 2024) first extracts file/method names from an issue report. Then, based on the names, it searches for corresponding classes and methods. It is an iterative search process and may continue to search for the called methods in the repository based on LLM feedback. AutoCodeRover may also be integrated with SBFL to refine localization and guide patch generation. However, it does not explicitly rank faulty methods; instead, it selects a method from the retrieved context and generates a patch.
In contrast to prior techniques, LLM4FL uses a structured, graph-based approach to analyze the code, enabling the LLM to systematically navigate the repository by considering caller-callee relationships. This eliminates LLM hallucination when searching and retrieving methods in the repository. Additionally, LLM4FL uses a divide-and-conquer strategy to mitigate token limitations by breaking down large methods into smaller, manageable groups, ensuring comprehensive analysis. Finally, LLM4FL refines fault localization results through verbal reinforcement learning, iteratively re-ranking suspicious methods for improved accuracy. We believe these contributions collectively enhance the accuracy and scalability of LLM-based fault localization.
3. Methodology
Traditionally, developers rely on dynamic execution data to facilitate fault localization: (i) code coverage is used for narrowing down executed code that changes the behavior incorrectly, whereas (ii) stack trace provides developers with insight into the program execution when the exception gets thrown, helping them diagnose the issue more effectively. However, directly applying Large Language Models (LLMs) for fault localization is infeasible. This is due to the large size of software artifacts, which frequently exceed token limitations, leading to input truncation and degraded performance  (Levy et al., 2024; Liu et al., 2024; Wu et al., 2023). To mitigate this issue, we propose LLM4FL, an LLM-agent-based fault localization system, which adopts a divide-and-conquer algorithm, which first divides the large software artifacts, and employs multiple LLM-Agents to localize faults autonomously.
Figure 1 provides an overview of our technique. LLM4FL consists of three LLM-Agents by using novel prompting techniques: (i) Context Extraction Agent, (ii) Debugger Agent, and (iii) Reviewer Agent to localize the fault iteratively. Specifically, the Context Extraction Agent utilizes a order-aware division and failure-reason guided prioritization prompting technique, which consists of two phases. In the division phase, a toolchain divides the large-scale code coverage data into small groups to fit within the LLM’s token limits. In the prioritization phase, the agent iteratively analyzes the divided code coverage, leveraging failed test cases and stack traces to identify potentially faulty methods. The Debugger Agent then performs graph-based retrieval-augmented code navigation prompting to locate code artifacts further to enhance the accuracy of fault ranking. Finally, the Reviewer Agent re-ranks the buggy method through verbal reinforcement learning prompting with the reviewer agent, ensuring more precise fault localization. Below, we provide more details on each of these phases.
Figure 1. Overview of LLM4FL for Multi-Agent Fault Localization, illustrating how agents collaborate to analyze software artifacts, extract failure reasoning, perform graph-based retrieval-augmented code navigation, and rank faulty methods using verbal reinforcement learning.
3.1. Context Extraction Agent
This agent defines a order-aware division and failure-reason guided prioritization prompting technique to divide code-coverage and iteratively prioritize faulty methods using stack trace and test code within each division. We describe the technique below.
3.1.1. Order-Aware Division Phase
To mitigate token size limitation, our LLM4FL first performs a division phase, where it first runs the test case to extract the list of method-level code coverage using GZoltar (Campos et al., 2012), which is denoted as $C$, containing a sequence of pairs ($m$, $s$), where $m$ denotes the method and $s$ denotes the set of statements within method $m$. We then divide $C$ into $K$ group of sequences of $C_{1},C_{2},\ldots,C_{k}$, where $K=\left\lceil\frac{\text{Input Token Length}}{\text{Token Limitation}}\right\rceil$, and each subset satisfies $|C_{i}|\leq\lfloor\text{Token Limitation}\rfloor$ to ensure the data fits within the LLM’s context window. For example, given an Input Token Length of 500K tokens, we divide this by the Token Limitation for GPT-4o-mini, which has a limit of 128K tokens (OpenAI, 2024). This results in $K=4$, each with 128K token limit.
Traditionally, the order in which divisions occur in divide-and-conquer algorithms does not impact the outcome. However, prior studies have shown that LLM may improve performance when the order of instructions is carefully considered  (Chen et al., 2024a). Inspired by this, we propose order-aware division, where we use a Spectrum-Based Fault Localization (SBFL) technique to sort $C$ based on their likelihood of being faulty. Specifically, we use Ochiai ranking to order $C_{1},C_{2},\ldots,C_{k}$, which is an efficient and unsupervised technique that assigns higher suspiciousness scores to statements that are executed more frequently by failing test cases and less frequently by passing ones (Abreu et al., 2006; Lou et al., 2021; Li et al., 2021; Cui et al., 2020; Wen et al., 2019; Qian et al., 2021).
3.1.2. Failure-Reason Guided Prioritization Phase
To prioritize faulty methods from the divided code coverage, one approach is to provide the agent with the stack trace and test case. However, prior studies have shown that incorporating summarized description can improve accuracy (Stiennon et al., 2020; Roit et al., 2023). Thus, we propose failure-reason-guided fault prioritization, which follows two steps. First, the input tokens are summarized into a failure-reason representation, capturing the test purpose, expected output, and failure reason. Second, the divided code coverage is prioritized based on the reason for failure. Below, we discuss each step in detail.
Figure 2. Context-Extraction Agent uses tool-chains to preprocess software artifacts for Lang-5 to emphasize the test failure context. For (i) stack-trace the agent prunes external libraries and (ii) test code, the agent prunes statements in the test code after the assertion failure.
Generate failure-reason from code artifacts. Our failure-reason generation makes the following three observations: (i) Firstly, the stack traces are verbose and include calls to external libraries unrelated to the fault. (ii) Secondly, some statements within a test case are irrelevant to the failure, specifically statements after the first assertion failures. (iii) Finally, the test case may call other helper methods that trigger the test failure (Peng et al., 2022). Hence, as shown in Figure 2, to mitigate (i), the Context Extraction Agent uses PreprocessTrace tool-chain to prune external execution in the stack trace. To mitigate (ii) and (iii), Context Extraction Agent uses PreprocessTest tool-chain, which uses static analysis to build an interprocedural call graph to extract all helper test methods called by the test and prunes all the statements appearing in the test after the failure point.
Given the preprocessed stack trace and test code, we prompt Context-Extraction Agents to generate failure-reason by answering test purpose, expected output, and failure reason. Below, we show a simplified example of the prompt:
“Analyze the test body and stack trace to generate a test failure reason summarizing the issue with expected output, and potential faulty location.”
## Input: ‘
{Test Code, Stack Trace}
## Response: ‘
{Failure-Reason}
An example of failure-reason generated for Lang-5 given stack trace and test code is:
## Test Purpose:
Test whether LocaleUtils.toLocale parses ’_GB’ as a locale string, expecting it to return a Locale object with an empty language and ’GB’ as the country.
## Failure Reason:
The actual output is an IllegalArgumentException due to an invalid locale format, as LocaleUtils.toLocale requires a language before the underscore. For example, ’_GB’ does not conform to the expected locale format. Update the test input to ’en_GB’ or modifying LocaleUtils.toLocale to handle ’_GB’ as a valid case. d
Failure-reason guided prioritization. Given Failure-reason, the Context Extraction Agent analyzes the covered methods in each group, previously denoted as $\{C_{1},C_{2},\dots,C_{K}\}$, to prioritize methods that are most related to the failure. Formally, prioritizing each set $C_{i}$ results in a prioritized subset $C^{\prime}_{i}$, where: $C^{\prime}_{i}\subseteq C_{i},\forall i\in\{1,2,\dots,K\}$. The prioritized methods from all subsets are then combined into a final prioritized set: $C^{\prime}=\bigcup_{i=1}^{K}C^{\prime}_{i}$. This final prioritized list $C^{\prime}$ is just a union of the ordered subsets, meaning no additional processing is introduced. This union process is necessary as LLMs are highly sensitive to initial order (Chen et al., 2024a), yet analyzing entire code coverage poses a token limitation. By dividing coverage into groups, prioritizing each, and unionizing subsets, we extract important context that are necessary for the Debugger Agent for fault analysis and localization in the next step.
3.2. Debugger Agent
Previously, the Context Extraction Agent prioritized $C$ to subsets $C^{\prime}$ to mitigate the LLM’s token limitation when analyzing large coverage data. While this technique can adapt to any LLM and token input size, prioritization may miss methods in the coverage that are important for fault localization. Inspired by how developers analyze the call dependencies to understand program execution and identify faults, we propose a graph-based retrieval-augmented code navigation strategy (Graph-RAG). This approach enables the Debugger Agent to utilize call graphs and method bodies for effective repository navigation, implementation analysis, generation of failure reasoning, and ranking the methods based on their likelihood of causing errors. We will detail this technique below.
3.2.1. Generating failure reasoning through call-graph-aware retrieval-augmented (Graph-RAG) code navigation.
The agent first uses code-coverage $C$ to construct the inter-procedural call-graph database $G=(V,E)$, where $V$ is a set of methods in the call-graph, and $E$ is a set of edges representing the caller-callee relationship. This graph is used by the tool-chains, get_MethodBody and get_CallGraph to facilitate agent-driven navigation within the call graph. LLM4FL then uses the prompt template below to drive the agent:
“Given (i) failure reasoning, (ii) stack trace, and (iii) test code, analyze, navigate, and enhance the failure reasoning for the given methods to identify faults. If additional implementation details are needed, extract the call graph and retrieve the caller or callee’s implementation. At the end, you should output the (i) all of the analyzed method, (ii) enhanced failure reasoning, (iii) fault ranking.”
## Input:
{Test Code, Stack Trace, Failure Reason, PrioritizedMethods}
## Response:
{Analyzed Method, Enhanced Failure Reason, Fault Ranking}
The agent begins the analysis of methods selected sequentially from the prioritized $C^{\prime}$. For each method, the agent retrieves its body using get_MethodBody and analyzes its implementation. It then queries get_CallGraph to identify relevant callers or callees, deciding whether to retrieve their implementations to enhance its understanding of code behavior. Throughout the iterative process of retrieval and analysis, every method examined, regardless of whether it was prioritized or invoked by a prioritized method, is assigned a failure reason by the agent. This information is then saved in a set referred to as $R$. This failure reasoning is a natural language explanation of each method’s failure analyzed in the code navigation process. These results are stored as pairs in a set: $R=\{(m_{i},r_{i})\mid m_{i}\text{ is analyzed},\;i=1,\dots,n\}$. Here, $n$ denotes the number of methods for which failure reasoning was generated.
Ranking methods based on failure reasoning results. Finally, given the failure reasoning $R$, the agent assigns each method $m_{i}\in\mathcal{R}$ an ordinal rank ($r^{*}_{i}$). Formally, the final ranked set $R^{*}$ is sorted in descending order based on the ordinal ranks:
$R^{*}=\text{sort}\left(\{(m_{i},r_{i},r^{*}_{i})\mid m_{i}\in R\}\right)$
where $r*_{i}$ denotes the ordinal rank that reflects each method’s likelihood of containing the fault, and $r_{i}$ explains the reasoning behind the method being faulty. The final ranked list $R^{*}$ is output in JSON format for subsequent analysis.
3.3. Reviewer Agent
Code review is important in software development to ensure software quality. Beyond traditional debugging, code review can help refine failure reasoning and improve fault localization. Inspired by this, we propose Reviewer Agent, which uses a novel prompting technique called Re-ranking through Verbal Reinforcement Learning to emulate a rigorous code-review process to refine fault localization. We describe the technique below.
3.3.1. Re-ranking through Verbal Reinforcement Learning
To refine the fault rankings $R^{*}$, Reviewer Agent adopts a verbal reinforcement learning process inspired by the Reflexion framework (Shinn et al., 2024). The agent iteratively critiques its rank trajectory by identifying inconsistencies, retrieving missing execution insights, and updating method prioritization to better reflect the failure’s root cause. After the iterations, it applies chain-of-thought to update the final ranking and return the fault localization results.
“Given (i) failure reasoning, (ii) stack trace, and (iii) test code, analyze and critique the initial ranking of suspicious methods. Reflect on the reasoning quality and identify missing or superfluous details. If additional insights are needed, extract the call graph and retrieve the caller or callee’s implementation to refine the ranking. At the end, you should output (i) the revised set of analyzed methods, (ii) the improved failure reasoning, and (iii) the final fault ranking with a possible fix.”
## Evaluating the Initial Ranking:
{Test Code, Stack Trace, Ranked List with Reasonings}
## Self-Critique and Refinement:
- Identify missing insights and remove unnecessary information.
- Retrieve method details using ‘get_MethodBody‘ if needed.
- Extract the call graph with ‘get_CallGraph‘ to enhance fault reasoning.
## Final Output:
{Analyzed Methods, Enhanced Failure Reasoning, Fault Ranking}
Evaluating and self-critiquing the initial ranking. The Reviewer Agent uses the initial ranking, $R^{*}$, provided by the Debugger Agent as the baseline trajectory for refinement. The agent enters a self-reflection phase, guided by Reflexion’s Self-Reflection Model ($M_{sr}$), to refine its ranking decisions through verbal reinforcement learning (Shinn et al., 2024). In this phase, the agent examines whether the initial ranking aligns with observed fault rankings by comparing the failure reasoning $r$, with caller-callee interactions and test coverage data, thereby identifying discrepancies such as misordered rankings or missing call dependencies. The agent generates natural language critiques and retrieves additional execution details if needed using tools such as get_MethodBody (to obtain complete method implementations) and get_CallGraph (to retrieve caller-callee relationships). The refined context is then incorporated in the Reviewer Agent’s iterative adjustment process.
Ranking adjustment through trajectory optimization. After the initial self-critique phase, the Reviewer Agent refines its initial ranking by updating each method’s rank $r^{*}$ and revising its failure reasoning $r$. Conceptualized as a trajectory optimization problem within the Reflexion framework, the agent, acting as an Actor ($M_{a}$), leverages reinforcement cues from its Self-Reflection Model ($M_{sr}$) to systematically re-order methods based on the updated evidence of inter-method interactions and failure reasoning. Each iteration integrates new feedback to adjust the ranking until it stabilizes or a preset iteration limit is reached.
Finalizing the ranking through chain-of-thought. After the iterative ranking adjustment, the Reviewer Agent finalizes the ranking one last time using chain-of-thought (Wang et al., 2024). To facilitate the thinking process, we ask the Reviewer Agent to generate a probable fix for every method in the final ranking by considering the updated ranking score $R^{*}$ and refined failure reasoning $R$. After generating all probable fixes, the agent then revisits all the information to do a final re-ranking. Formally, the final ranked list $R^{*}_{final}=\text{sort}\left(\{(m_{i},r^{*}_{i},r_{i},f_{i})\mid m_{i}\in C^{\prime}\}\right)$, where $f_{i}$ denotes the fix generated for method $m_{i}$. The final ranked list $R$, encoded in JSON format, provides a structured, machine-readable prioritization along with human-interpretable failure justifications.
4. STUDY DESIGN AND RESULTS
In this section, we first describe the study design and setup. Then, we present the answers to the research questions.
Table 2. An overview of our studied projects from Defects4J v2.0.0. #Faults, LOC, and #Tests show the number of faults, lines of code, and tests in each system. Fault-triggering Tests shows the number of failing tests that trigger the fault.
Project #Faults LOC #Tests Fault-triggering Tests
Cli 39 4K 94 66
Closure 174 90K 7,911 545
Codec 18 7K 206 43
Collections 4 65K 1,286 4
Compress 47 9K 73 72
Csv 16 2K 54 24
Gson 18 14K 720 34
JacksonCore 26 22K 206 53
JacksonXml 6 9K 138 12
Jsoup 93 8K 139 144
Lang 64 22K 2,291 121
Math 106 85K 4,378 176
Mockito 38 11K 1,379 118
Time 26 28K 4,041 74
Total 675 380K 24,302 1,486
Benchmark Dataset. To answer the RQs, we conducted the experiment on 675 faults across 14 projects from the Defects4J benchmark (V2.0.0) (Just et al., 2014). Defects4J provides a controlled environment to reproduce faults collected from various types and sizes of projects. Defects4J is widely used in prior automated fault localization research (Lou et al., 2021; Sohn and Yoo, 2017; Chen et al., 2022; Zhang et al., 2017). We excluded three projects, JacksonDatabind, JxPath, and Chart, from Defects4J in our study because we encountered many execution errors and were not able to collect test coverage information for them. Table 2 gives detailed information on the projects and faults we use in our study. The faults have over 1.4K fault-triggering tests (i.e., failing tests that cover the fault). The sizes of the studied projects range from 2K to 90K lines of code. Note that since a fault may have multiple fault-triggering tests, there are more fault-triggering tests than faults.
Evaluation Metrics. According to prior findings, debugging faults at the class level lacks precision for effective location (Kochhar et al., 2016). Alternatively, pinpointing them at the statement level might be overly detailed, omitting important context (Parnin and Orso, 2011). Hence, in keeping with prior work (Benton et al., 2020; B. Le et al., 2016; Li et al., 2019; Lou et al., 2021; Vancsics et al., 2021), we perform our fault localization process at the method level. We apply the following commonly-used metrics for evaluation:
Recall at Top-N. The Top-N metric measures the number of faults with at least one faulty program element (in this paper, methods) ranked in the top N. The result from LLM4FL is a ranked list based on the suspiciousness score. Prior research (Parnin and Orso, 2011) indicates that developers typically only scrutinize a limited number of top-ranked faulty elements. Therefore, our study focuses on Top-N, where N is set to 1, 3, 5, and 10.
Implementation and Environment. To collect test coverage data and compute results for baseline techniques, we utilized Gzoltar (Campos et al., 2012), an automated tool that executes tests and gathers coverage information. For the LLM-based components, we employed OpenAI’s gpt-4o-mini-2024-07-18, a more cost-effective yet capable LLM (OpenAI, 2024). We used LangChain to develop LLM4FL (Langchain, 2024). We designed the prompts to be concise to minimize token usage and to allow more room for analysis-related information and code. To reduce the variations in the output, we set the temperature parameter to 0 during model inference.
RQ1: How does LLM4FL perform compared with other fault localization techniques?
Motivation. In this RQ we evaluate the fault localization accuracy of LLM4FL by comparing with various baseline techniques.
Approach. We compare LLM4FL’s fault localization accuracy with five baselines representing different methodological families: Ochiai (statistical) (Abreu et al., 2006), DeepFL (deep neural network) (Li et al., 2019), Grace (graph neural network) (Lou et al., 2021), DepGraph (graph neural network) (Rafi et al., 2024), AutoFL (LLM) (Kang et al., 2024), and AgentFL/SoapFL (Qin et al., 2024) (LLM).
Ochiai (Abreu et al., 2006) is a widely recognized statistical fault localization technique known for its high efficiency, making it a common baseline for comparison (Lou et al., 2021; Li et al., 2021; Cui et al., 2020; Wen et al., 2019; Qian et al., 2021; Rafi et al., 2024). As such, we use Ochiai to rank the methods during the segmentation process and include it as a baseline for accuracy comparison.
DeepFL (Li et al., 2019) is a deep-learning-based fault localization technique that integrates spectrum-based and other metrics such as code complexity, and textual similarity features to locate faults. It utilizes a Multi-layer Perceptron (MLP) model to analyze these varied feature dimensions. We follow the study (Li et al., 2019) to implement DeepFL and include the SBFL scores from 34 techniques, code complexity, and textual similarities as part of the features for the deep learning model. Grace (Lou et al., 2021) utilized graph neural networks (GNN). It represents code as a graph and uses a gated graph neural network to rank the faulty methods. DepGraph (Rafi et al., 2024) is another GNN-based technique that further improves Grace by enhancing code representation in a graph using interprocedural call graph analysis for graph pruning and integrating historical code change information.
AutoFL (Kang et al., 2024) is an LLM-based fault localization approach that provides the LLM with a failing test and method descriptions to gather relevant coverage data, then repeats the process to assign inverse scores and rank candidate methods by averaging results across runs. SoapFL (also known as AgentFL(Qin et al., 2024) organizes fault localization into structured phases, including test failure comprehension, code navigation, and iterative LLM-based scoring, leveraging enhanced method documentation to refine suspicious methods. While both AutoFL and SoapFL used OpenAI’s GPT-3.5 for their experiments, for our evaluation, we adapted both techniques to use the same LLM version as LLM4FL (i.e., gpt-4o-mini-2024-07-18) for comparison.
Table 3. Fault localization accuracy (Top-1, 3, 5, and 10) for 675 faults from Defect4J V2.0.0. The numbers in the parenthesis show the percentage difference compared to LLM4FL.
Techniques Top-1 Top-3 Top-5 Top-10
Ochiai 121 (169.42%) 260 (63.46%) 340 (39.41%) 413 (20.10%)
DeepFL 257 (26.85%) 353 (20.40%) 427 (11.01%) 468 (5.98%)
Grace 298 (9.40%) 416 (2.16%) 486 (-2.47%) 541 (-8.32%)
DepGraph 359 (-9.19%) 481 (-11.64%) 541 (-12.38%) 597 (-16.92%)
AutoFL 275 (18.55%) 393 (8.14%) 423 (12.06%) 457 (8.53%)
SoapFL/AgentFL 311 (4.82%) 414 (2.66%) 455 (4.18%) 478 (3.77%)
LLM4FL 326 425 474 496
Results. LLM4FL outperforms the LLM-based baselines, AutoFL and AgentFL, by achieving a 185.55% and 4.82% improvement in Pass@1, respectively. Tables 3 show the fault localization results of LLM4FL and the baseline techniques. Among the three LLM-based techniques, LLM4FL achieves a better Top@N across all values of N. In the Top-1 metric, LLM4FL locates the correct fault in 326 cases, compared to AutoFL’s 275 and AgentFL’s 311, representing an 18.55% and 4.82% improvement, respectively. Similarly, in Top-3, Top-5, and Top-10, LLM4FL an improvement between 8.14% to 12.06% over AutoFL and 2.66% to 4.18% for AgentFL. These numbers highlight LLM4FL’s ability to pinpoint faulty methods more accurately. For cost, we compute the total dollars spent by multiplying each approach’s token usage by the per-million-token price (e.g., \$0.150 for input, \$0.600 for output). SOAPFL averages about \$0.055 per bug, AutoFL averages about \$0.065, while LLM4FL is at around \$0.050 per bug making our approach comparably cost-effective.
LLM4FL shows higher Top-1 and Top-3 compared to most other non-LLM-based techniques. For the Top-1 metric, LLM4FL scores 326, which is 169.42% higher than Ochia’s score of 121, 26.84% higher than DeepFL’s score of 257, and 9.39% better than Grace’s result of 298. One exception is DepGraph, which achieves a Top-1 of 359, 8.64% higher than LLM4FL. As the range expands to Top-3 and beyond, LLM4FL demonstrates its robustness, significantly outperforming DeepFL and maintaining strong performance alongside Grace. LLM-based techniques have several advantages over traditional techniques such as DepGraph. First, techniques like LLM4FL leverage pre-trained LLMs in a zero-shot setting, which can be easily applied to systems with insufficient training data. Second, techniques like DepGraph require additional model training that can take days. LLM-based techniques leverage generic pre-trained LLMs without specific fine-tuning or re-training. Finally, LLM-based techniques can explain the decision, which helps with adaption (Kang et al., 2024; Qin et al., 2024). Hence, LLM-based techniques have a strong potential to enhance fault localization by offering greater adaptability, reducing training overhead, and providing interpretable explanations.
LLM4FL achieves a higher Top-1 compared to other LLM-based techniques, AutoFL and AgentFL, by 18.55% and 4.82%, respectively. It also has competitive results compared to supervised techniques like DeepFL and Grace by leveraging a pre-trained LLM in a zero-shot setting.
RQ2: How do different components in LLM4FL affect the fault localization accuracy?
Motivation. LLM4FL employs a number of components, and each of these components plays a distinct role in the overall process. Hence, in this RQ, we conduct an ablation study by removing each component separately and study their impact on fault localization accuracy. The findings may inspire future studies on adapting the components for similar tasks.
Approach. To evaluate the impact of each component, we designed four different configurations:
LLM4FLw/o CodeNav removes the code navigation mechanism, which essentially means the Debugger agent no longer does fault navigation by retrieving the call graphs. Instead, the LLM4FL uses a single prompt to perform fault localization without fault navigation. This configuration tests whether using the caller-callee information improves the ranking and selection of faulty methods or if a single round of analysis is sufficient.
LLM4FLw/o Division removes the order-aware division of the covered methods, giving the agents the entire coverage data at once instead of dividing it into smaller, manageable groups. Coverage segmentation addresses token limitations in LLMs, so removing it explores the impact of feeding the full dataset to the agents in one step. We aim to see how handling large amounts of data in a single input influences the fault localization result, as it may overwhelm the model or reduce precision.
LLM4FLw/o Reflexion removes the verbal reinforcement learning technique, which is used to allow agents to review and refine their initial ranking. Without this step, the agents rely solely on their initial assessments without iterative improvements.
Table 4. Impacts of removing different components in LLM4FL on Top-1, 3, 5, and 10. The numbers in the parentheses show the percentage changes compared to LLM4FL with all the components.
Techniques Top-1 Top-3 Top-5 Top-10
LLM4FL 327 425 473 494
LLM4FLw/o CodeNav 273 (-16.51%) 378 (-11.06%) 409 (-13.53%) 409 (-17.21%)
LLM4FLw/o Division 251 (-23.24%) 341 (-19.76%) 365 (-22.83%) 381 (-22.87%)
LLM4FLw/o Reflexion 290 (-11.31%) 400 (-5.88%) 436 (-7.82%) 459 (-7.09%)
Results. While all components help improve the results, including coverage division and code navigation provide the largest improvement to fault localization results (23% and 17% in Top-1). Table 4 shows the Top-1, 3, 5, and 10 scores when each component is removed. Removing coverage divisions has the largest overall impact across all scores, reducing Top-N by 19% to 23%. Removing prompt chaining has the second largest impact (11% to 17%). At the individual project level, these two components also have the largest impact in Top-1 in 9/13 studied projects. Our finding shows that employing sorted coverage grouping following the divide and conquer technique and agent communication significantly improves fault localization results. Future research should consider these techniques when designing fault localization techniques.
Although there is no oracle during the fault localization process, asking LLMs to self-reflect still helps improve the overall Top-1 by 11%. LLMs often suffer from hallucinations, especially when there is a lack of feedback from external oracles (Xu et al., 2024; Huang et al., 2023). Even though we did not provide any groundtruth or external feedback to LLM, we found that reflexion is still effective in improving fault localization results. We speculate that verbal reinforcement learning helps the model improve its results by revisiting the suspicious methods it ranked earlier, creating a feedback loop. In this process, we analyze the results for each group of methods and combine these group-specific results into a final ranking. The model can then spot mistakes or gaps in its logic, leading to better results. Self-reflection brings 6% to 11% improvement across the Top-N metrics. Our finding suggests that future studies should consider self-reflection even if there is no external feedback. Our finding highlights the effectiveness of self-reflection, which should be considered in future fault localization results.
The results show that each component of LLM4FL contributes to its overall fault localization performance, with coverage division and code navigation having the largest positive impact. Removing these components leads to significant declines in accuracy, confirming their role in finding faults in a large code-base.
RQ3: Does order matter in the initial list of methods provided to the LLM?
Motivation. LLM4FL divides the coverage data into different divisions using an order-aware division strategy to address the token size limitation of LLMs. We sort the methods using the Ochiai scores before the division, though different sorting mechanisms may affect the final fault localization result. Although LLM4FL eventually visits and assesses every method, a recent study (Chen et al., 2024a) observes that the order of premises affects LLM’s results. However, whether this effect extends to software engineering tasks, particularly fault localization, remains unclear. Hence, in this research question, we investigate whether the order of methods within the groups affects the LLM’s fault localization performance.
Approach. To test the effect of method ordering, we experiment with three distinct sorting strategies: LLM4FLExecution, LLM4FLOchiai (the default sorting in LLM4FL), and LLM4FLDepGraph to sort the methods before the divide-and-conquer step.
LLM4FLExecution: We use the unsorted list of methods executed during testing, as generated by Gzoltar (Campos et al., 2012). This default list represents the natural execution order of the methods, with no explicit ranking or prioritization. By providing the LLM with methods based on the execution sequence, we establish a control case to measure its performance without any ranking influence.
LLM4FLOchiai: As discussed in Section 3.1, we apply Ochiai to sort the methods before the divide-and-conquer process. Ochiai is unsupervised and is efficient to compute. We hypothesize that providing the LLM with methods sorted by their suspiciousness score will lead to more effective fault localization, as the model will focus on the most likely faulty candidates earlier in the process.
LLM4FLDepGraph: This approach uses the ranking produced by DepGraph, a state-of-the-art Graph Neural Network (GNN)-based fault localization technique (Li et al., 2015; Rafi et al., 2024), to sort the methods. DepGraph ranks methods based on structural code dependencies and code change history. As shown in RQ1, DepGraph shows the highest fault localization accuracy among all the techniques, surpassing LLM4FLOchiai. By examining the fault localization result after sorting the methods using DepGraph’s scores, we can better study if the initial order affects LLM’s results, even though LLM eventually visits all the methods.
Figure 3. Fault localization results when using different method sorting strategies during the segmentation process.
Results. Method ordering has a significant impact on LLM’s fault localization result, with up to 22% difference in Top-1 (from 299 to 366). Figure 3 shows the fault localization results when using different sorting strategies. When methods were presented in the execution order, LLM4FLExecution achieved a Top-1 score of 299, 402 for Top-3, 431 for Top-5, and 462 for Top-10. This performance establishes a baseline, showing how the LLM behaves without strategic ordering. However, sorting methods with the lightweight Ochiai scores resulted in noticeable improvements across all Top-N, where LLM4FLOchiai improved the Top-1 score to 323, an 8% increase over LLM4FLExecution.
LLM4FLDepGraph provides further improvement to the already-promising result of DepGraph, indicating method ordering is critical to LLM4FL, or LLM-based fault localization in general. LLM4FLDepGraph achieved the highest Top-1 score of 366, which significantly outperforms both LLM4FLExecution and LLM4FLOchiai. The improvement was consistent across all the metrics. We also see that LLM4FLDepGraph has better Top-1, 3, and 5 scores compared to DepGraph. This consistent improvement underscores the importance of method ordering in enhancing the accuracy of LLM-based fault localization. Namely, if the initial order is closer to the group truth, the final localization result also tends to be more accurate.
Our finding establishes a new research direction for LLM-based fault localization, or any software engineering tasks that take a list of software artifacts as input. Future studies may study how different premises of ordering affect other software engineering tasks, and how to combine traditional software engineering techniques to pre-process LLM’s input to further improve the results.
The initial method ordering significantly impacts the accuracy of LLM-based fault localization, with Top-1 scores varying by up to 22%. Future research should explore different ordering strategies and how traditional software engineering techniques can be integrated to optimize LLM performance further.
5. Threats to Validity
Internal Validity. A potential threat to internal validity is the risk of data leakage in large language models (LLMs), where the model might have been exposed to the benchmark data during training. Nevertheless, Ramos et al. (Ramos et al., 2024) found that newer and larger models trained on larger datasets exhibit limited evidence of leakage for defect benchmarks. Since we utilize GPT-4o-mini, a large model trained using a tremendous amount of data, it likely shares similar characteristics in reducing memorization risks. Additionally, following prior work, we ensure no content related to the project name, human-written bug report, or bug ID is entered into ChatGPT to minimize the risk of data memorization (Qin et al., 2024).
External Validity. The primary threat to external validity is the generalizability of our results. Our evaluation is based on Defects4J, a well-established dataset in the software engineering community. Although this dataset includes real-world bugs, the systems studied are primarily Java-based. Future studies may extend our study to other programming languages or domains.
Construct Validity. Construct validity relates to whether the metrics we used accurately measure the performance of fault localization techniques. We used widely accepted Top-N metrics commonly utilized in prior fault localization studies. However, our results are based on the assumption that developers primarily focus on the top-ranked faulty methods. Although this assumption aligns with previous research, different development practices could influence the effectiveness of our approach.
6. Conclusion
In this paper, we introduced LLM4FL, an LLM-agent-based fault localization approach. It utilizes multiple specialized LLM agents, including Context, Debugger, and Reviewer, to iteratively refine and improve the accuracy of fault localization through order-aware prioritization, graph-based code navigation, and self-reflection using verbal reinforcement learning. Evaluated on the Defects4J (V2.0.0) benchmark, LLM4FL demonstrated significant improvements over existing approaches, achieving an 18.55% increase in Top-1 accuracy compared to AutoFL and 4.82% improvement over SoapFL. Further enhancements, including coverage segmentation and iterative refinement, increased accuracy by up to 22%. Future work will explore expanding LLM4FL’s capabilities for larger and more diverse codebases, further refining the agent collaboration and reasoning mechanisms.